# EasyOcc: 3D Pseudo-Label Supervision for Fully Self-Supervised Semantic Occupancy Prediction Models

### v3

@@ Abstract
Self-supervised models have recently achieved notable advancements, particularly in the domain of semantic occupancy prediction. These models utilize sophisticated loss computation strategies to compensate for the absence of ground-truth labels. For instance, techniques such as novel view synthesis, cross-view rendering, and depth estimation have been explored to address the issue of semantic and depth ambiguity. However, such techniques typically incur high computational costs and memory usage during the training stage, especially in the case of novel view synthesis. To mitigate these issues, we propose 3D pseudo-ground-truth labels generated by the foundation models Grounded-SAM and Metric3Dv2, and harness temporal information for label densification. Our 3D pseudo-labels can be easily integrated into existing models, which yields substantial performance improvements, with mIoU increasing by 45%, from 9.73 to 14.09, when implemented into the OccNeRF model. This stands in contrast to earlier advancements in the field, which are often not readily transferable to other architectures. Additionally, we propose a streamlined model, EasyOcc, achieving 13.86 mIoU. This model conducts learning solely from our labels, avoiding complex rendering strategies mentioned previously. Furthermore, our method enables models to attain state-of-the-art performance when evaluated on the full scene without applying the camera mask, with EasyOcc achieving 7.71 mIoU, outperforming the previous best model by 31%. These findings highlight the critical importance of foundation models, temporal context, and the choice of loss computation space in self-supervised learning for comprehensive scene understanding. The code and dataset will be made available open-source at https://github.com/seamie6/EasyOcc
Figure 1: Overview of the EasyOcc Model Architecture: Image features are first extracted from the six surrounding camera views using a 2D image encoder. These features are then projected into a 3D feature volume via bilinear sampling. The resulting voxel grid is processed by a 3D CNN to generate the final semantic occupancy prediction. Loss is computed by comparing the predicted output (top) with the corresponding 3D pseudo-labels (bottom).
@@ I Introduction
Recent progress in machine learning has resulted in highly capable models for tackling complex tasks, particularly Vision Language Models (VLM) such as Grounding DINO [grounding_dino], and Visual Foundation Models (VFM) including Metric3Dv2 [metric3dv2] and Segment Anything (SAM) [sam]. Trained on large datasets with hundreds of millions of parameters, these models enable capabilities such as zero-shot depth estimation and semantic segmentation via Metric3Dv2 and Grounded-SAM [grounded_sam, grounding_dino, sam] respectively. Each model holds promise for autonomous perception, especially in self-supervised semantic occupancy prediction, which involves assigning semantic labels to voxels in a discretized 3D grid without manually annotated ground-truth labels for supervision. The lack of ground-truth labels leads to the issue of semantic and depth ambiguity, which foundation models help address. Existing methods leverage foundation models to generate 2D pseudo-labels for supervision: SelfOcc employs OpenSeeD [openseed] for semantic supervision, while both OccNeRF and GaussianOcc utilize Grounded-SAM. GaussTR [gausstr] integrates CLIP [clip] for feature supervision and Metric3Dv2 for depth supervision. More recently, the generation of 3D pseudo-labels has been studied [ago, autoocc], and our work aims to advance this research further by integrating 3D pseudo-labels into existing models and striving for holistic scene representation.
Despite the deployment of such strong foundation models, a key challenge that persists in this domain is the complexity of training strategies. SelfOcc uses novel view synthesis and a multi-layer perceptron (MLP) to predict signed distance functions for occupancy and semantics. OccNeRF applies NeRF-based volume rendering [nerf] for depth and semantics, while GaussianOcc uses a two-stage process: estimating transformations between subsequent sample coordinate frames, and then applying Gaussian Splatting [gaussian_splatting] for rendering. These methods reflect a trend toward extracting maximal information from camera images using increasingly complex techniques. For these models, all loss is computed in 2D camera space, as pseudo-ground-truth labels are generated from the camera images. Representing these labels in 3D eliminates the need for the costly 2D rendering of the scene, allowing direct scene representation and loss computation in 3D.
@@
We propose a method that uses Grounded-SAM and Metric3Dv2 to generate 3D pseudo-labels for direct loss computation in 3D space, which easily allows the model to understand semantics and spatial geometry in one loss function, pseudo-loss. To increase label density, we aggregate temporal samples, limited to static objects to avoid duplicating dynamic objects, a known issue in self-supervised [selfocc, occnerf, gaussianocc] and weakly supervised models [renderocc, gsrender]. Prior work [sparseocc, ttgaussocc, stcocc] has shown the value of temporal information for improving performance. Aggregating temporal information is typically performed at inference time via the aggregation features, thereby introducing additional computational overhead during model deployment. Our method avoids this issue. In summary, 3D pseudo-labels provide several important advantages. Firstly, they remove the requirement for novel view synthesis and depth estimation during training, thereby lowering computational demands. Secondly, they enable efficient aggregation of temporal information, which plays a crucial role in spatial understanding. Lastly, they allow models to develop a more comprehensive understanding of the scene, leading to improved performance not only in regions visible to the camera but, more importantly, in areas occluded from the current viewpoint.
@@
Furthermore, our 3D pseudo-labels can be easily added to existing models as an auxiliary loss to boost performance. We explore the integration of these labels in three previous models: SelfOcc, OccNeRF, and GaussianOcc. As recent advances often develop in isolation, we present a complementary method to enhance model compatibility and generalization. Additionally, we introduce EasyOcc, a streamlined model that solely uses our generated 3D pseudo-labels for loss computation, demonstrating that complex rendering techniques are not necessary for noteworthy model performance. Unlike earlier self-supervised methods, our approach aligns more with supervised frameworks [daocc, occfusion], with a simpler architecture, which is illustrated in Figure 1. EasyOcc requires no LiDAR supervision or deployment of foundation models at inference.
@@
  • 3D Pseudo-Labels: We introduce an approach that leverages Grounded-SAM and Metric3Dv2 to generate 3D pseudo-labels for loss computation directly in 3D space.
  • Seamless Integration: Our labels can be effortlessly integrated into existing models via an auxiliary loss function, yielding improvements of 43% in mIoU.
  • Segmentation of Dynamic Classes: Accurate segmentation of dynamic classes is critical for autonomous perception. In the case of SelfOcc, incorporating our 3D pseudo-labels improves prediction segmentation performance by 627%.
  • Holistic Scene Representation: The proposed 3D pseudo-labels enable a more comprehensive representation of the scene, resulting in a 219% increase in mIoU for OccNeRF when evaluated across the entire voxel grid.
This paper is structured as follows: Section II reviews prior work on semantic occupancy prediction models and pseudo-labels. Section III details the generation of our 3D pseudo-labels and provides an analysis of their quality. Following this, Section IV details the EasyOcc model and 3D pseudo-label integration into existing architectures. Section V compares models incorporating our 3D pseudo-labels with state-of-the-art (SOTA) models, highlighting quantitative and qualitative improvements through ablation studies. Finally, Section VI presents concluding remarks.
@@ II Literature Review
This section is structured as follows: Subsection II-A reviews semantic occupancy prediction models and analyzes the three models selected for modification, along with other SOTA models we shall compare against. Subsection II-B discusses pseudo-labels used in this space.
II-A Semantic Occupancy Prediction
@@
In autonomous perception, Bird’s Eye View (BEV) methods have historically been dominant due to their simple yet effective scene representation [simplebev, liftsplat, bevformer, bevfusion]. Recently, semantic occupancy prediction has gained attention, driven by benchmark datasets [surroundocc, occ3d, nucraft, openocc] with accurate annotations, generated from manually labeled nuScenes LiDAR data [nuScenes]. This shift led to the creation of notable supervised semantic occupancy prediction models, with improvements from techniques such as Gaussian Splatting, multi-modal fusion, and object deduplication [gaussian_splatting, daocc, occfusion, renderocc, gsrender]. Following this, self-supervised counterparts of these models emerged, particularly due to their flexibility in training strategy. In this study, we modify three self-supervised models: SelfOcc, OccNeRF, and GaussianOcc, with each model’s pipeline discussed in this section, along with other SOTA models that we will be comparing against.
@@
SelfOcc employs an RN-50 image encoder followed by a 3D encoder to generate a 3D scene representation, inspired by BEVFormer [bevformer]. An MLP then predicts signed distance field (SDF) values, color, and semantic features from the 3D volume, for rendering depth, color, and semantics [selfocc]. Rendered depth supports multi-frame photometric consistency, rendered color is compared to the camera image, while semantics are compared against 2D pseudo-labels from OpenSeeD. Semantics and occupancy are both computed via the SDF, with both contributing to the final scene representation. In our implementation, pseudo-loss is applied exclusively to the semantic voxel, with the occupancy voxel excluded from both the pseudo-loss computation and the final scene representation, as detailed in Subsection IV-B. Our loss computation technique aligns with that of supervised models, as both compare a predicted voxel grid against voxel labels in 3D space.
OccNeRF follows a similar pipeline, using an RN-101 for feature extraction and bilinear sampling to project image features into 3D voxel space. These features are refined using a 3D CNN [occnerf]. Bilinear sampling outperforms depth-based splatting methods seen in Lift-Splat-Shoot [liftsplat], though multi-scale deformable attention provides superior performance at higher computational cost [simplebev]. Following the 3D CNN, NeRF rendering is employed to render both depth and semantic information, with depth supporting multi-frame photometric consistency, while semantics are compared against pseudo-labels from Grounded-SAM [grounded_sam, grounding_dino, sam]. Similar to SelfOcc, pseudo-loss is computed on the predicted 3D voxel of semantic logits.
@@
GaussianOcc builds on OccNeRF with Gaussian rasterization [gaussian_splatting] for rendering both semantics and depth [gaussianocc]. For use in multi-frame consistency, it estimates pose transformations using a 6D pose network instead of ground-truth poses, which is more effective, given the nuScenes dataset’s lack of $z$-axis translation in ego-vehicle transformations [nuScenes]. Pseudo-loss is computed in the same manner as in OccNeRF.
GaussTR represents the scene using 3D Gaussians, which are refined through self-attention and image cross-attention mechanisms [gausstr]. The model utilizes CLIP and Metric3Dv2 for pseudo-supervision. Similarly, TT-OccCamera also models the scene with 3D Gaussians, however, it is a test-time compute method, which eliminates the need for pretraining. GaussianFlowOcc refines Gaussians using image cross-attention, induced self-attention, and induced temporal attention in relation to previous Gaussians [gaussianflowocc]. Additionally, it incorporates temporal 2D pseudo-labels to enhance performance. The model further leverages Grounded-SAM and Metric3Dv2 for pseudo-supervision.
AutoOcc addresses semantic occupancy prediction using a VLM-only framework [autoocc]. It begins by computing attention maps from surround-view images and candidate object categories [InternVL, lin2024trainingfreeopenendedobjectdetection, abnar2020quantifyingattentionflowtransformers]. These attention maps are then input into VLMs to extract semantic and depth features [unidepth, fastersam, fastsam], which are subsequently projected into 3D space. The label refinement process involves flow estimation and smoothing of instance-level predictions using 3D Gaussian Splatting. AutoOcc presents their work as both a prediction paradigm but also a pseudo-labelling method. In contrast, our approach investigates the integration of our 3D pseudo-labels into additional models, enabling distillation into significantly smaller architectures. AutoOcc does not explore this direction.
@@
Our proposed model, EasyOcc, shares a similar pipeline with GaussianOcc. However, it omits several components: there is no pose estimation, novel view synthesis using Gaussian Splatting, or multi-frame photometric consistency. Instead, EasyOcc uses a single loss function, pseudo-loss, comparing voxel predictions directly with our generated 3D pseudo-labels. Despite its simplified design, EasyOcc outperforms other models that employ complex techniques in the mIoU metric, as detailed in the results Section V.
We did not pursue implementing 3D pseudo-labels in GaussTR, as its Gaussian-based scene representation would require significant adjustments to maintain practical training times due to the inefficiencies of the Gaussian-to-voxel splatting module. TT-OccCamera, GaussianFlowOcc, and AutoOcc are currently not open-source and hence omitted for modification.
Figure 2: Our Method of Generating 3D Pseudo-Labels: Grounded-SAM labels are first projected into 3D space using Metric3Dv2 depth maps and camera pose information to produce a semantic point cloud. To densify the point cloud, we aggregate 13 temporal samples while filtering out dynamic objects to avoid duplication. The resulting densified point cloud is then passed to a voxelization module to generate 3D pseudo-voxel labels.
@@ II-B Pseudo-Labels
The use of 2D pseudo-labels in self-supervised semantic occupancy models has been extensively studied, particularly through the application of VLMs [grounded_sam, clip, talk2dino] and VFMs [unidepth, metric3dv2]. These models have demonstrated significant utility across various domains, including medical applications [MedSAM] and other autonomous perception tasks [seamie_paper]. They play a crucial role in addressing the challenge of missing ground-truth labels, particularly in resolving semantic and depth ambiguities. VLMs help mitigate semantic ambiguity by leveraging both spatial and linguistic cues to produce pixel-level semantic maps. Depth ambiguity, while partially addressed using multi-frame photometric consistency, is significantly reduced through the use of metric depth VFMs [metric3dv2, unidepth, depthpro], which provide accurate pixel-level depth maps, shown to increase model performance substantially [gausstr, gaussianflowocc]. Semantic maps and depth maps serve as supervision signals against renders of the semantic voxel grid. Nonetheless, a noticeable performance gap persists between supervised and self-supervised models, indicating a need for a more nuanced integration of foundation models, potentially by focusing on the dimensionality of the labels.
@@
More recently, the use of 3D pseudo-labels has gained attention, which aligns with the space in which final predictions are made. A notable example is AGO [ago], which combines Grounded-SAM [grounded_sam] with LiDAR point cloud data, utilizing multi-frame aggregation, point cloud ray casting, and semantic voting to generate richer labels for training. However, this approach necessitates equipping the vehicle with a LiDAR sensor, which introduces significant cost and complexity, including the need for careful synchronization with the camera system. Additionally, the output generated by the aforementioned AutoOcc can be considered a form of 3D pseudo-labels.
Our approach to 3D pseudo-label generation does not rely on LiDAR data and instead combines Metric3Dv2 [metric3dv2] with Grounded-SAM [grounded_sam, grounding_dino, sam] to produce 3D pseudo-ground-truth labels. Metric3Dv2 is selected for its reliable depth estimation across different camera views, while Grounded-SAM is utilized for its high-quality semantic segmentation. The generated labels undergo further refinement through techniques such as outlier removal, occupancy thresholding, and temporal aggregation. These 3D pseudo-labels enable models to learn more effective scene representations, particularly when evaluated over the full voxel grid, a setting that has not been thoroughly investigated in prior work.
III Methodology I: 3D Pseudo-Labels
This section begins with the generation process for our 3D pseudo-labels in Subsection III-A. Following this, subsection III-B evaluates the quality of our 3D pseudo-labels when aggregating multiple samples and altering the occupancy threshold value. A comparison against ground truth data is also provided.
III-A Generation of 3D Pseudo-Labels

### v4

# Easy3D-Labels: Supervising Semantic Occupancy Estimation with 3D Pseudo-Labels for Automotive Perception

@@ Abstract
In perception for automated vehicles, safety is critical not only for the driver but also for other agents in the scene, particularly vulnerable road users such as pedestrians and cyclists. Previous representation methods, such as Bird’s Eye View, collapse vertical information, leading to ambiguity in 3D object localisation and limiting accurate understanding of the environment for downstream tasks such as motion planning and scene forecasting. In contrast, semantic occupancy provides a full 3D representation of the surroundings, addressing these limitations. Furthermore, self-supervised semantic occupancy has seen increased attention in the automated vehicle domain. Unlike supervised methods that rely on manually annotated data, these approaches use 2D pseudo-labels, improving scalability by reducing the need for labour-intensive annotation. Consequently, such models employ techniques such as novel view synthesis, cross-view rendering, and depth estimation to allow for model supervision against the 2D labels. However, such approaches often incur high computational and memory costs during training, especially for novel view synthesis. To address these issues, we propose Easy3D-Labels, which are 3D pseudo-ground-truth labels generated using Grounded-SAM and Metric3Dv2, with temporal aggregation for densification, permitting supervision directly in 3D space. Easy3D-Labels can be readily integrated into existing models to provide model supervision, yielding substantial performance gains, with mIoU increasing by 45% and RayIoU by 49% when applied to OccNeRF on the Occ3D-nuScenes dataset. Additionally, we introduce EasyOcc, a streamlined model trained solely on these 3D pseudo-labels, avoiding the need for complex rendering strategies, and achieving 15.7 mIoU on Occ3D-nuScenes. Easy3D-Labels improve scene understanding by reducing object duplication and enhancing depth estimation accuracy, as reflected by improvements in the RayIoU metric. These findings highlight the importance of foundation models, temporal information, and 3D loss formulation in self-supervised learning for comprehensive scene understanding. Our Easy3D-Labels are available open-source on Mendeley
Figure 1: Easy3D-Labels generation: We project semantic labels into 3D using depth maps and employ temporal aggregation and object filtering for enhanced label quality.
@@ I INTRODUCTION
Safety in the domain of automated vehicle perception is critical to establish trust for both drivers and other road users [40]. In 2021, the global annual road traffic deaths were estimated at 1.19 million, with pedestrians accounting for 21% of these fatalities [50]. It is evident that vulnerable road users account for a large proportion of road injuries and deaths. This emphasises the need for accurate and timely detection of agents in the scene to enable the vehicle to predict trajectories and make decisions in high-risk situations.
Detection accuracy is closely tied to how the scene is represented, as restrictive representations can limit both perceptual coverage [30] and the completeness of spatial reasoning [32]. Following recent progress in machine learning, many advanced approaches for automated vehicle perception have been developed [52, 12]. One prominent approach is semantic occupancy estimation [49], whose discretised 3D representation enables more flexible scene modelling compared to earlier methods such as Bird’s Eye View and 3D bounding boxes that contain the aforementioned limitations.
Specifically, self-supervised semantic occupancy estimation is especially advantageous, largely due to its reduced reliance on manually annotated occupancy labels, offering improved scalability compared to supervised approaches [17]. However, these models still depend on labels in the 2D image space, using Vision Language Models (VLM) [27, 36] and Visual Foundation Models (VFM) [22, 14] to address semantic and depth ambiguities in the absence of ground-truth 3D labels. To enable supervision in 2D, models typically use novel view synthesis techniques to render the 3D scene in image space, employing NeRF-based volume rendering [33, 55] or 3D Gaussian Splatting [21, 9, 19]. However, relying on 2D rendering introduces high computational cost and leads to limitations such as bias towards nearby objects [12], inaccurate detections, and object duplication [41].
@@
To address these issues, we propose a 3D labelling technique, Easy3D-Labels, which generates labels by projecting Grounded-SAM [37] 2D semantics into 3D space using Metric3Dv2 [14] depth maps, illustrated in Figure 1. This allows the model to jointly learn semantics and spatial geometry through a single pseudo-loss function. Furthermore, we aggregate temporal samples, limited to static objects, to avoid duplicating dynamic objects to increase density. Prior work has shown the value of temporal information for improving performance [44, 57, 25]. The aggregation of temporal data in the supervision pipeline reduces additional computational overhead, as it is typically performed at inference time. Easy3D-Labels provide several important advantages. Firstly, the removal of the need for novel view synthesis and depth estimation during training reduces computational cost. Secondly, they enable efficient aggregation of temporal information, which is important for spatial understanding. Lastly, they support more comprehensive scene understanding in both visible and occluded regions
@@
Furthermore, Easy3D-Labels can be easily added to existing models as an auxiliary loss to boost performance. We explore the integration of these labels in three previous models: SelfOcc, OccNeRF, and GaussianOcc. As recent advances often develop in isolation, we present a complementary method to enhance model compatibility and generalization. Additionally, we introduce EasyOcc, a streamlined model that solely uses Easy3D-Labels for loss computation, demonstrating that complex rendering techniques are not necessary for noteworthy model performance. Furthermore, our model requires no LiDAR supervision or deployment of foundation models at inference.
@@
  • Easy3D-Labels: We introduce an approach that leverages Grounded-SAM and Metric3Dv2 to generate 3D pseudo-labels for loss computation directly in 3D space. Our 3D pseudo-labels can be effortlessly integrated into existing models via an auxiliary loss function, yielding improvements of 45% in mIoU.
  • Segmentation of Dynamic Classes: Accurate segmentation of dynamic classes is critical for safety in autonomous perception. In the case of SelfOcc, incorporating our 3D pseudo-labels improves pedestrian segmentation performance by over 600%.
  • Holistic Scene Representation: Our labels enable a more comprehensive scene representation, resulting in a 49% increase in RayIoU for OccNeRF.
This paper is structured as follows: Section II reviews related work, Sections III and IV describe Easy3D-Labels and the EasyOcc model, Section V presents results and ablations, and finally Section VI concludes the paper.
@@ II LITERATURE REVIEW
In Subsection II-A, we review semantic occupancy models, and in Subsection II-B, we discuss prior use of 3D pseudo-labels.
II-A SEMANTIC OCCUPANCY ESTIMATION
@@
In perception for automated vehicles, Bird’s Eye View (BEV) methods have historically been dominant due to their simple yet effective scene representation [11, 34, 24, 28]. Recently, semantic occupancy estimation has gained attention, driven by benchmark datasets [49, 45, 64, 46] with accurate annotations, generated from manually labeled LiDAR data from the nuScenes automated vehicles dataset [6]. This shift led to the creation of supervised occupancy estimation models [43, 38], with improvements from techniques such as Gaussian Splatting [18, 16, 59, 39], multi-modal fusion [13, 53, 31], and object deduplication [41]. Following this, self-supervised counterparts of these models emerged, particularly due to their flexibility in training strategy, requiring no manually annotated ground truth labels. In this study, we modify three self-supervised models, SelfOcc, OccNeRF, and GaussianOcc, which all follow a common pipeline of 2D image encoding, 3D feature lifting, and 3D voxel refinement prior to rendering.
@@
SelfOcc employs an MLP to predict signed distance field (SDF) values, color, and semantic features from the 3D volume, for rendering depth, color, and semantics [17]. Depth supports multi-frame photometric consistency, color for comparison to the RGB image, and semantics against 2D pseudo-labels from OpenSeeD [58]. Semantics and occupancy are both computed via the SDF, with both contributing to the final scene representation. In our implementation, pseudo-loss is applied exclusively to the semantic voxel.
@@
OccNeRF deploys NeRF-style volume rendering to render both depth and semantic information, with depth supporting multi-frame photometric consistency, while semantics are compared against pseudo-labels from Grounded-SAM [37]. GaussianOcc builds on OccNeRF with Gaussian rasterization [21] for rendering both semantics and depth [9]. For use in multi-frame consistency, it estimates pose transformations using a 6D pose network instead of ground-truth poses, which is more effective, given the nuScenes dataset’s lack of $z$-axis translation in ego-vehicle transformations [6]. For both methods, pseudo-loss is applied to the semantic voxel grid.
Other state-of-the-art methods employ Gaussian scene representations, which prove beneficial for reduced memory consumption due to their sparse nature. Methods employ techniques including self-attention and image cross-attention mechanisms [19, 5], foundation models during inference [19, 57], and temporal flow modelling [5], with supervision labels originating again from foundation models [47, 14, 37, 2, 20]. AutoOcc is an estimation and labelling pipeline which uses numerous attention mechanisms [8, 26, 1] and foundation models [35, 54, 60] in a test-time manner. However, this method does not explore the deployment of its labels into existing models for enhancement.
@@
Our proposed model, EasyOcc, shares a similar pipeline with GaussianOcc. However, it omits several components: pose estimation, novel view synthesis, and multi-frame photometric consistency. Instead, EasyOcc solely leverages Easy3D-Labels for supervision. Despite its simplified design, EasyOcc outperforms other models that employ complex training paradigms in the mIoU and RayIoU, as detailed in Section V.
@@ II-B PSEUDO-LABELS
The use of 2D pseudo-labels in self-supervised semantic occupancy models has been extensively studied, particularly through the application of VLMs [37, 36, 2] and VFMs [35, 14]. These models have demonstrated utility across various domains, including medical applications [56, 29, 51, 10], and robotics [15, 7]. In perception for automated vehicles, they can play a crucial role in addressing the challenge of missing ground-truth labels, particularly in resolving semantic and depth ambiguities. VLMs help mitigate semantic ambiguity by leveraging both spatial and linguistic cues to produce pixel-level semantic maps. Depth ambiguity, while partially addressed using multi-frame photometric consistency, is significantly reduced through the use of metric depth VFMs [14, 35, 4], which provide accurate pixel-level depth maps, shown to increase model performance substantially [19, 5]. Semantic maps and depth maps serve as supervision signals against renders of the semantic voxel grid. Nonetheless, a noticeable performance gap persists between supervised and self-supervised models, indicating a need for a more nuanced integration of foundation models, potentially by focusing on the dimensionality of the labels.
@@
More recently, the use of 3D pseudo-labels has gained attention, which aligns with the space in which final estimations are made. A notable example is AGO [23], which combines Grounded-SAM [37] with LiDAR point cloud data, utilizing multi-frame aggregation, point cloud ray casting, and semantic voting to generate richer labels for training. However, this approach necessitates equipping the vehicle with a LiDAR sensor, which introduces high cost and complexity, including the need for careful synchronization with the camera system. Additionally, the output generated by the aforementioned AutoOcc can be considered a form of 3D pseudo-labels.
Our proposed method, Easy3D-Labels, does not rely on LiDAR and instead combines Metric3Dv2 [14] for depth estimation with Grounded-SAM [37] for semantic segmentation to generate 3D pseudo-ground-truth labels. These labels are further refined through outlier removal, occupancy thresholding, and temporal aggregation, enabling models to learn a more holistic scene representation.
III METHODOLOGY I: EASY3D-LABELS
This section outlines the generation of Easy3D-Labels in Subsection III-A, followed by an evaluation of their quality in Subsection III-B, including comparisons with ground truth.
III-A GENERATION OF EASY3D-LABELS